OpenStudy (anonymous):

x^2+y^2=26 y-x=-6 Solving system of non linear equations Please Help!?!?!?!?!!!!

OpenStudy (nikvist):

\[x^2+y^2=26\] \[y-x=-6\quad\Rightarrow\quad y=x-6\] \[x^2+(x-6)^2=26\,,\,x^2+x^2-12x+36=26\,,\,2(x^2-6x+5)=0\] \[2(x-1)(x-5)=0\quad\Rightarrow\quad (x,y)=\{(1,-5),(5,-1)\}\]